OpenStudy (anonymous):

The equation below can be used to find the cost, including 8% tax, of p pounds of potatoes. Julie wants to know how many pounds of potatoes she can buy for $11.34. Solve the equation and write your answer. 1.75p(1.08) = 11.34 Show your work.

OpenStudy (anonymous):

6 pounds

OpenStudy (anonymous):

1.08*1.75p=11.34 1.89p =11.34 p = 11.34/1.89 p = 6 got any question throw them at me

OpenStudy (anonymous):

thanku

OpenStudy (anonymous):

no problem